OpenStudy (anonymous):

Solve for x. State any restrictions on the variable. a(3tx-2b)=c(dx-2)

OpenStudy (anonymous):

\[\Large a(3tx-2b)=c(dx-2)\] \[\Large a3tx-2ab=cdx-2c\] \[\Large a3tx-cdx=2ab-2c\] \[\Large x(a3t-cd)=2ab-2c\] \[\Large x=\frac{2ab-2c}{a3t-cd}\]